w= yh-5yc^3 ( which is 5yc to the 3rd power), for y

y=
i was told no solution from someone in here-y= ?? there is an answer

Answer by richwmiller(17219)

w= yh-5yc^3
w= y(h-5c^3)
w/(h-5c^3)=y
y=w/(h-5c^3)